OpenStudy (anonymous):

solve the equation: (3x-2/12) - (1/6) = (1/6)

OpenStudy (anonymous):

(3x-2/12) - (1/6) = (1/6) (3x-2/12)= (2/6) 3x= 3/6 x=1/6

OpenStudy (anonymous):

wait where did the -2/12 go?

OpenStudy (anonymous):

I changed it into 1/6 and added it to the other side

OpenStudy (anonymous):

oh. Okay

OpenStudy (anonymous):

wait i think i put that wrong. the (3x-2)/12. would it still be the same. because the 3x-2 is all together.

OpenStudy (anonymous):

It would be different...Type up the actual problem?

OpenStudy (anonymous):

\[((3x-2)/12) - (1/6) = (1/6)\]

OpenStudy (anonymous):

((3x−2)/12)−(1/6)=(1/6) ((3x−2)/12)=2/6 3x-2=4 3x=6 x=2

OpenStudy (anonymous):

Thanks (:

OpenStudy (anonymous):

np